A copper vessel of mass 1000g contains 500g of water at 50C. An ice cube of mass 100g at temperature 30C is dropped into the vessel.

a)Does the ice melt? Explain in a few words how you reached to your answer.

b) What is the final temperature?

c) Repeat part b if at the same time as dropping in the ice cube, wealso drop in a 200g piece of glass at a temperature of 110C.

Specific heats in cal/g.C

Ice=0.5 water=1.0 Copper=0.1 Glass=0.2

Latent heat of water: Vapourization =540cal/g Fusion=80cal/g
Answer #1

A)

yes the ice will melt , as the initial temperture of water and vessel is high

and the ice will melt to come to an equilibrium temperature.

B)

let the final temperature is Tf

0.10 * 1000 * (50 - Tf) + 500 * 1 *(50 - Tf) = 80 * 100 + 30 * 100 * 0.50 + 100 * 1 * Tf

Tf= 29.3 degree

the final temperature is 29.3 degree

C)

for also dropping the glass ,

0.10 * 1000 * (50 - Tf) + 500 * 1 *(50 - Tf) + 200 * 0.20 *(110 - Tf) = 80 * 100 + 30 * 100 * 0.50 + 100 * 1 * Tf


solving for Tf

Tf = 33.65 degree C

the final temperature is 33.65 degree C